area of a circle is 220 cm2.Find area of square inscribed in it.

Grade:12th pass

13 Answers

vannala shivanand
8 years ago
area of the circle is 220 means  3.14*r2 . then area o fthe square is area of the circle divided by 3.14. that means area of the square is 70.
anish
8 years ago
the answer is 140 cm2. From the area of the circle will get the value of r.after that find the value of the diameter of the circle i.e 2r.But  if we see for the square  the diameter is the length of the diagonal of the square i.e 2r=sq .root 2a(a= side of the square).and from here we will get a2=140 cm2.
sri tanish
8 years ago
area of circle is 220 => radius = 701/2   => diameter = 280½        diagonal of square =  >  area = [280{1/2}2]/2  =>  140
 
Bodhoyita Nath
7 years ago
Area of a circle = 220=> Πr² = 220=> 22/7 r² = 220=> 1/7 r² = 10=> r² = 10×7=> r² = 70=> r = √70.We know, diameter = 2×radiusTherefore, diameter = 2×√70Now, area of square = 1/2× (diagonal) ² =1/2×2×√70×2×√70 =(√70)² × 2 = 70 × 2 = 140.Therefore, the area of square is 140cm².
test
7 years ago
area of circle is 220 => radius = 701/2   => diameter = 280½        diagonal of square =  >  area = [280{1/2}2]/2  =>  140
test
7 years ago
area of circle is 220 => radius = 701/2   => diameter = 280½        diagonal of square =  >  area = [280{1/2}2]/2  =>  140..
Gowtham
6 years ago
area of a circle = 220
π r2 = 220
r = 220/π 
r = 70.
diameter = 2r
d= (2r)2
d = 4 * r2
d = 4 * 70 = 280
area of circle = ½ * d = 280/2 = 140 .
 
 
Probin pegu
6 years ago
Area of a circle = 220=> Πr² = 220=> 22/7 r² = 220=> 1/7 r² = 10=> r² = 10×7=> r² = 70=> r = √70. If we let the square is ABCD and the centre is O then from triangle AOB we get AB²= OA² + OB² =(√70)² +(√70)² = 70 +70 AB² =140 AB = √140Therefore the area of the square ABCD= AB× AB= √140 × √140= 140 cm²
Balamurugan
6 years ago
 = 280/2 = 140 . a2a= d2/2we know that in square, area  = 70. diameter d=2r  d2=(2r)2d2=4r2d2=4x70=280rπ = 220/r= 220  r πArea of a circle
ankit singh
3 years ago
Area of a circle = 220=> Πr² = 220=> 22/7 r² = 220=> 1/7 r² = 10=> r² = 10×7=> r² = 70=> r = √70.We know, diameter = 2×radiusTherefore, diameter = 2×√70Now, area of square = 1/2× (diagonal) ² =1/2×2×√70×2×√70 =(√70)² × 2 = 70 × 2 = 140.Therefore, the area of square is 140cm². 
π r2 = 220
r = 220/π 
r = 70.
diameter = 2r
d= (2r)2
d = 4 * r2
d = 4 * 70 = 280
area of circle = ½ * d = 280/2 = 140 .
Arjun Raj
3 years ago
Area of circle = 220 
πr=220
r2=220/π   put π=22/7
r2=70
r=√70
diameter, ,d = 2r = diagonal of square
d=2√70
Area of square=1/2×d2
                           =1/3 × 4 ×70
                          = 140
RISHIKA
3 years ago
DEAR FRIEND
Area of the circle = 220 cm. sq. The radius of the circle is half the diagonal of the square. Therefore, Area of the square whose vertices are on the circle= 140 cm².
THANK U
Joseph Antony
3 years ago
 
area of a circle = 220
π r2 = 220
r = 220/π 
r = 70.
diameter = 2r
d= (2r)2
d = 4 * r2
d = 4 * 70 = 280
area of circle = ½ * d = 280/2 = 140 .